danaw2001:

A 62.0 kg water skier at rest jumps from the dock into a 775 kg boat at rest on the east side of the dock. If the velocity of the skier is 4.50 m/s as she leaves the dock, what is the final velocity of the skier and boat?

DanielleLove:

62*4.5+775*0 = (62+775)*(-v) 279 = 837(-v) -837v = 279 v = -(279/837) v = -558